OpenStudy (anonymous):

find the average squared distance between the point of R = {(x,y): 0 <= x <= 4, 0 <= y <=5} and the point (4,5)

OpenStudy (anonymous):

\[\frac{ 1 }{ 20 }\int\limits_{0}^{5}\int\limits_{0}^{4}[(4-x)^{2}+(5-y)^{2}]dxdy\] figured it out